A rocket-powered sled accelerates from rest. After t=10s it has traveled a distance x=400m. What is its speed in km/h at this point?
1
Expert's answer
2019-07-15T09:13:28-0400

Hello .

v=x/t v=400/10=40m/c =144km/h  Answer:144km/h